Sergio invests part of his 90,000 at 4% and the rest at 3%. How much is invested at each rate if he makes a total income of 3,100 annually?
--HELP ME PLEASE I CANT DO IT BY MYSELF :(

Found 2 solutions by lynnlo, mananth:
Answer by lynnlo(4176)

4%=====================1 yr. value==========$46,833.39
total interest=========================$1,833.39
3%=====================1 yr. value==========$46,368.71
total interest=========================$1,368.71

Answer by mananth(16949)

Part I 3.00% per annum ------------- Amount invested =x
Part II 4.00% per annum ------------ Amount invested = y =
90000
Interest----- 3100

Part I 3.00% per annum ---x
Part II 4.00% per annum ---y
Total investment
x + 1 y= 90000 -------------1
Interest on both investments
3.00% x + 4.00% y= 3100
Multiply by 100
3 x + 4 y= 310000.00 --------2
Multiply (1) by -3
we get
-3 x -3 y= -270000.00
Add this to (2)
0 x 1 y= 40000
divide by 1
y = 40000
Part I 3.00% $ 50000
Part II 4.00% $ 40000

CHECK
50000 --------- 3.00% ------- 1500.00
40000 ------------- 4.00% ------- 1600.00
Total -------------------- 3100.00

m.ananth@hotmail.ca